A projectile is fired in such a way that its horizontal range is equal to 3 times its maximum height. What is the angle of projection?


1
Expert's answer
2021-05-18T11:09:49-0400

Write the equation that ties the velocity and angle of projection with the range:


"R=\frac{v^2\sin(2\theta)}{g}."

The height:


"H=\frac{v^2\sin^2\theta}{2g}."

We read that


"R=3H, \\\space\\
\frac{v^2\sin(2\theta)}{g}=\frac{3v^2\sin^2\theta}{2g},\\\space\\
2\cos\theta=\frac32\sin\theta,\\\space\\
\tan\theta=\frac43,\\\space\\
\theta=53°."